Skip to content
CalcGospel 國際數學圖譜
返回

S1 2023 June Q6

A Level / Edexcel / S1

IAL 2023 June Paper · Question 6

题目

Problem

Three events AA, BB and CC are such that

P(A)=0.1P(B∣A)=0.3P(A∪B)=0.25P(C)=0.5P(A)=0.1\qquad P(B\mid A)=0.3\qquad P(A\cup B)=0.25\qquad P(C)=0.5

Given that AA and CC are mutually exclusive

(a) find P(A∪C)P(A\cup C)

(1)

(b) Show that P(B)=0.18P(B)=0.18

(3)

Given also that BB and CC are independent,

(c) draw a Venn diagram to represent the events AA, BB and CC and the probabilities associated with each region.

(5)
题目中文翻译

三个事件 AA、BB 和 CC 满足题目所列的概率条件。

已知 AA 与 CC 互斥。

(a) 求 P(A∪C)P(A\cup C)。

(b) 证明 P(B)=0.18P(B)=0.18。

另知 BB 与 CC 相互独立。

(c) 绘制维恩图,表示事件 AA、BB、CC 以及各区域相应的概率。

解答

(a)

解法一

思路

展开

AA 和 CC mutually exclusive,所以交集为 00,联合概率直接相加。

答题过程

展开 P(A∪C)=P(A)+P(C)=0.1+0.5=0.6.\begin{align*} P(A\cup C)=P(A)+P(C)=0.1+0.5=0.6. \end{align*}

(b)

解法一

思路

展开

先由条件概率求 P(A∩B)P(A\cap B),再代入加法公式求 P(B)P(B)。

答题过程

展开

Using

P(B∣A)=P(A∩B)P(A),\begin{align*} P(B\mid A)=\frac{P(A\cap B)}{P(A)}, \end{align*}

we get

P(A∩B)=0.3(0.1)=0.03.\begin{align*} P(A\cap B)=0.3(0.1)=0.03. \end{align*}

Now

P(A∪B)=P(A)+P(B)−P(A∩B).\begin{align*} P(A\cup B)=P(A)+P(B)-P(A\cap B). \end{align*}

So

0.25= 0.1+P(B)−0.03P(B)= 0.18.\begin{align*} 0.25=&\,0.1+P(B)-0.03\\ P(B)=&\,0.18. \end{align*}

(c)

解法一

思路

展开

先放 A∩B=0.03A\cap B=0.03。因为 AA 与 CC 互斥,所以 A∩C=0A\cap C=0,三者交集也为 00。再用 BB 与 CC 独立求 B∩CB\cap C。

答题过程

展开

Since BB and CC are independent,

P(B∩C)=P(B)P(C)=0.18(0.5)=0.09.\begin{align*} P(B\cap C)=P(B)P(C)=0.18(0.5)=0.09. \end{align*}

Also,

P(A∩B)=0.03.\begin{align*} P(A\cap B)=0.03. \end{align*}

Since AA and CC are mutually exclusive,

P(A∩C)=0,P(A∩B∩C)=0.\begin{align*} P(A\cap C)=0,\qquad P(A\cap B\cap C)=0. \end{align*}

The AA only region is

0.10−0.03=0.07.\begin{align*} 0.10-0.03=0.07. \end{align*}

The BB only region is

0.18−0.03−0.09=0.06.\begin{align*} 0.18-0.03-0.09=0.06. \end{align*}

The CC only region is

0.50−0.09=0.41.\begin{align*} 0.50-0.09=0.41. \end{align*}

The outside region is

1−(0.07+0.03+0.06+0.09+0.41)= 0.34.\begin{align*} 1&-(0.07+0.03+0.06+0.09+0.41)\\ =&\,0.34. \end{align*}

So the Venn diagram should contain the regions

A only=0.07,A∩B=0.03,B only=0.06,B∩C=0.09,C only=0.41,outside=0.34,\begin{gathered} A\text{ only}=0.07,\quad A\cap B=0.03,\quad B\text{ only}=0.06,\\ B\cap C=0.09,\quad C\text{ only}=0.41,\quad \text{outside}=0.34, \end{gathered}

with the A∩CA\cap C regions equal to 00.